OpenStudy (anonymous):

18. Write an equation of a line whose graph is parallel to the graph of y = 3x – 10.

OpenStudy (texaschic101):

first we need to find the slope of the equation given. We do this because a parallel line will have the same slope. Do you know how to find the slope of y = 3x - 10 ??

OpenStudy (mkapad01):

y=mx+b m is the slope, any equation with the same slope is parallel to the prior line

OpenStudy (texaschic101):

not always true...because an equation can have the same slope, but it cannot have the same y intercept, or it will be consistent...the same line. So any equation that has a slope of -2 and any y intercept number except 10.

OpenStudy (texaschic101):

so y = -2x + 3 or y = -2x + 4 are parallel. But y = -2x + 10 is not parallel because the y intercepts would be the same

OpenStudy (mkapad01):

we know it's not the same exact line, so we know it's not going to have the same y-intercept

OpenStudy (texaschic101):

oops....slope is 3x, not -2...I do not know where I got that from...brain fart

OpenStudy (texaschic101):

so a parallel line would be : y = 3x + 4, or y = 3x + 5....it cannot be y = 3x - 10

OpenStudy (anonymous):

I still dont understand

OpenStudy (texaschic101):

or.. In y = mx + b form, the slope is the m position and the y intercept is the b position y = 3x = 10 slope(m) = 3 (y intercept) = (0,10)..x = 0 and y = 10 now we sub y = mx + b 10 = 3(0) + b 10 = b parallel equation is : y = 3x + 10

OpenStudy (texaschic101):

typo...y = 3x - 10

OpenStudy (anonymous):

ok thank you

OpenStudy (texaschic101):

that is the original equation.....the parallel one would be y = 3x + 10

OpenStudy (anonymous):

you mean y=3x-10

OpenStudy (texaschic101):

y = 3x - 10 is the original equation....a parallel equation to this would be y = 3x + 10

OpenStudy (mkapad01):

I disagree, any number would fit for b

OpenStudy (texaschic101):

true...I am not saying this is the only parallel equation. Any number would fit for b except -10

OpenStudy (anonymous):

y=3x=10 i thougt it was y=3x-10